# Interpreting the Win Ratio in Hierarchical Composite Endpoints: Challenges, Limitations, and Perspectives with Examples from Chronic Kidney Disease Trials

### v1

@@ Abstract
Win statistics based methods have gained traction as a method for analyzing Hierarchical Composite Endpoints (HCEs) in randomized clinical trials, particularly in cardiovascular and kidney disease research. HCEs offer several key advantages, including increased statistical power, the mitigation of competing risks, and hierarchical weighting of clinical importance for different outcome components. While, as summary measures, the win ratio (WR) along with the Net Benefit (NB) and the Win Odds (WO) provide a structured approach to analyzing HCEs, several concerns regarding their interpretability remain. In this paper, we explore critical aspects of WR interpretation that have received limited attention. Specifically, we discuss the challenge of defining an appropriate estimand in the context of HCEs using the WR, the difficulties in formulating a relevant causal question underlying the WR, and the dependency of the WR on the variance of its components, which complicates its role as an effect measure. Additionally, we highlight the non-collapsibility of the WR, akin to hazard and odds ratios, further complicating its interpretation. While the WR remains a valuable tool in clinical trials, its inherent limitations must be acknowledged to ensure its appropriate application and interpretation in regulatory and clinical decision-making.
Keywords: Win Ratio, Hierarchical Composite Endpoints, Win Statistics, Estimand, Causal Interpretation, Collapsibility.
@@ 1 Introduction
In recent years, there has been a notable surge in the analysis of Hierarchical Composite Endpoints (HCE) using the Win Ratio (WR) in randomized clinical trials. Mao’s study[1], based on a search of ClinicalTrials.gov, highlights a significant rise in the application of WR analyses. Within the realm of cardiovascular and kidney disease, several major outcome trials have employed the WR in some capacity to compare a novel treatment intervention with a comparator.[2, 3, 4, 5, 6, 7, 8] In this article, we focus on two-armed clinical trials, specifically evaluating the efficacy of an active treatment against a comparator."
@@ 1.1 Hierarchical composite endpoint
A common feature of the HCEs used in these trials is that they typically involve one terminal event, one or more non-terminal events, and one or more continuous/ordinal outcomes. For example, Pocock et al.[2], defined a heart failure HCE as "...a hierarchical composite of death, number of HF events, time to first HF event, or a $\geq$5-point difference in Kansas City Cardiomyopathy Questionnaire (KCCQ) total symptom score change at 90 days." Similarly, Heerspink et al.[4] proposed and investigated a kidney disease HCE comprising "(1) all-cause mortality; (2) kidney replacement therapy defined as dialysis for at least 28 days or kidney transplantation; (3) sustained GFR <15 mL/min per 1.73 $\text{m}^{2}$ for at least 28 days; (4) sustained GFR decline from baseline for at least 28 days of 57%; (5) 50%; (6) 40%; or (7) GFR slope". GFR slope refers to the individual-level, annualized rate of decline in glomerular filtration rate (GFR), which represents kidney function.
@@
While the ordering of the HCE may seem straightforward, it is not always the case. To illustrate with an extreme example, suppose a trial where all subjects in the comparator arm systematically experience the terminal event slightly earlier than those in the active arm. At the same time, all active arm subjects experience a non-terminal severe event within the HCE, and none in the comparator arm do. In this case, win statistics based on the HCE will indicate superiority for the active arm due to the small benefit on the terminal event, yet the HCE ignores the increased frequency of non-terminal severe events in the active arm. These limitations are also present in other analyses of composite endpoints; for example, a time-to-first-event analysis of this type of data would encounter similar issues. To address these concerns, it is in general advisable to also evaluate the components individually.[9]
@@ 1.2 Win Ratio
The WR, as proposed by Pocock et al. (2012) [10], serves as a summary measure for analyzing HCEs when comparing two treatment arms. In the unmatched version of this approach, each subject from the novel treatment arm is compared to each subject from the comparator arm to determine a "winner" for each comparison. Initially, the analysis determines whether one subject wins over the other on the first HCE component. If a clear winner emerges, that subject is declared the winner for that specific subject comparison. If a winner cannot be determined, a tie is declared, and the process moves on to the next component. This sequence continues until a winner is declared for the entire comparison or until a tie is declared for the last component, resulting in an overall tie for that particular subject comparison. Let $P_{t}$ denote the proportion of pairs when treatment wins, $P_{c}$ denote the proportion of pairs when control wins, and $P_{\text{tie}}=1-P_{t}-P_{c}$ denote the proportion of ties. The WR is then estimated as $P_{t}/P_{c}$.
@@
Other often used summary measures for HCEs are the net benefit (NB) and the Win Odds (WO). The NB, also referred to as the Win Difference, computes the difference in win probabilities between treatment arms, i.e. $P_{t}-P_{c}$. The WO, estimated as $(P_{t}+0.5P_{\text{tie}})/(P_{c}+0.5P_{\text{tie}})$, adjusts for tied outcomes by allocating half of the ties to the wins for each arm, and thus ’penalizing’ trials with many ties. While these three win statistics differ in their interpretations and sensitivity to ties, their corresponding statistical tests are asymptotically equivalent under standard regularity conditions.[11, 12] We will focus on the WR in this article, but briefly mention NB and WO when relevant.
When applying the WR for analyzing HCE with a time-to-event component in a trial without a fixed follow-up for all subjects, there is no assurance of transitivity; if subject A wins against subject B and subject B wins against subject C, there is no guarantee that subject A will win against subject C. This situation can arise when comparing subjects observed over varying time intervals, leading to comparisons relying on overlapping observation intervals; for more details, refer to Oakes (2016) [13], and Oakes (2025).[14]
@@
All of the benefits and issues presented in this section are to be considered when applying a WR to HCEs. However, in this article, we will direct our focus toward different aspects related to the interpretation of the WR estimate, which we believe have received little to no attention so far. First, we will delve into the challenge of defining a suitable estimand for which a WR analysis of an HCE would provide a proper estimator, particularly formulating the clinical question that the WR would aptly address. Second, we will tackle the difficulty of defining a relevant causal question that can be answered by a WR analysis. Third, we will address the fact that a WR on an HCE with one or more continuous components will be contingent on the variance of those components, thereby complicating the interpretation of the WR as an effect measure. As a consequence, the WR is also a non-collapsible measure (similar to hazard or odds ratios), further complicating its interpretation.
@@
All these factors collectively lead to the acknowledgment that viewing the WR as an effect measure is not straightforward.
@@ 2 Estimand
The estimand framework was introduced in 2020 with the publication of the ICH E9 addendum.[15] This framework provides a structured approach to defining estimands in clinical trials, ensuring clarity, transparency, and interpretability of clinical trial results, thereby improving decision-making processes in drug development and regulatory assessments. The key attributes of the estimand include:
@@
  • Treatment: Assign a novel or alternative treatment, on top of standard of care.
@@
  • Population-level Summary: The odds that a random subject assigned to the intervention has a better outcome than a random subject assigned to control.
@@
A fundamental consideration when defining an estimand for a WR-based analysis is its dependency on the maximum follow-up time and the censoring distribution. Since the WR is based on pairwise comparisons between subjects, its estimation is affected by the time window over which these comparisons are conducted.[13] To define a valid estimand that is independent of censoring, one approach is to restrict WR calculations to a fixed time horizon $t$, such as 24 months.[16, 17] This improves comparability across studies and reduces interpretational complexities from differential follow-up, as the contributions of the more severe endpoints comprising the HCE tend to increase with longer follow-up time; thus, the relative contributions of the different endpoints will generally vary between patients with shorter versus longer follow-up durations.
Alternatively, if clinically justified, the WR can be assumed to remain constant over time, allowing estimation of a single value representing the entire study period.[18, 19] Despite these considerations, hypothesis testing for the standard WR remains valid, even if its interpretation depends on follow-up duration. The usual statistical tests for WR, such as asymptotic methods or permutation-based approaches, remain robust for evaluating the null hypothesis of no treatment effect. However, careful interpretation of the WR estimate is necessary to ensure it aligns with the clinical question of interest.
@@
If the primary clinical interest lies in comparing the location of a continuous endpoint between treatment arms, in a population where a significant number of deaths is anticipated during the trial, one might consider employing an HCE with time to death as the first component, and the continuous endpoint as the second in order to circumvent interpretational challenges resulting from early termination of follow-up by "truncation by death". However, if we are to remain faithful to the strict estimand framework, the technical challenges of managing intercurrent events, such as death, in the comparison of continuous outcomes should not dictate the choice of the estimand. Specifically, if the estimand aims to assess the treatment effect on the continuous endpoint, the composite strategy of handling intercurrent events through an HCE will not adequately address the estimand of interest. Moreover, if applied to the continuous endpoint, the WR as a treatment effect measure does not estimate the difference in the locations of the continuous distributions (as can be seen from the below example of normally distributed random variables, WR is related to the standardized mean difference and not the mean difference); rather, it assesses the odds that a randomly selected subject assigned to the intervention has a better outcome than a randomly selected subject assigned to the control group.
On the other hand, in the CKD example the primary clinical interest lies in the treatment effect on kidney function decline. Thus endpoint components of interest would be GFR slope and relevant clinical events such as dialysis or kidney transplantation. Interpreting GFR values after dialysis or transplantation is not straightforward. Hence, the estimation of GFR slope in the presence of these intercurrent events, which are by themselves highly relevant for the evaluation of kidney function, may also be challenging. To overcome these issues, employing the HCE for the estimand as outlined in the CKD example above might be an attempt for a holistic assessment of kidney function.
@@ 2.1 Causal interpretation
As mentioned in section 2, the ICH E9 addendum on estimands emphasizes the importance of clearly defining the treatment effect of interest in clinical trials. The addendum encourages the use of causal language and frameworks to describe these effects, ensuring that the treatment effect is interpretable in a causal manner. For an introductory text on causal inference please refer to Hernán & Robins.[20]
@@
where comparisons are done within subjects. However this is not identifiable in a randomized experiment.[21, 22]
There is no guarantee that these two are similar. One can even construct examples with opposing conclusions. As an illustration, consult the so-called Hands paradox.[23, 21]
In Table 1, we present the factual and counterfactual responses for three subjects under two treatments. Here, Y(0) represents the response under the comparator treatment, while Y(1) denotes the response under the active treatment. A higher response value indicates a better outcome. In this toy example, the WR (population-level), comparing all outcomes in the two arms is: ((0+0+0) + (0+1+0) + (0+1+1)) / ((1+1+1) + (1+0+1) + (1+0+0)) = 3/6 = 0.5, whereas the $\widetilde{\text{WR}}$ (individual-level), doing the comparisons within individuals (0+1+1)/(1+0+0) = 2/1 = 2. Thus, not only are they different, but they are pointing in opposite directions; one would lead to a conclusion of a beneficial individual-level treatment effect, whereas the population-level would suggest a harmful treatment effect.
@@
The non-identifiable individual-level $\widetilde{\text{WR}}$ has a straightforward interpretation for patients and physicians: "what is the mean odds of obtaining a better outcome vs a worse outcome when being exposed to treatment 1 vs 0". However, the interpretation of the population-level WR for a subject/physician is more dubious: "what is the mean odds of a better outcome for a subject exposed to treatment 1 vs another subject exposed to treatment 0". While this population-level comparison is valid in a causal sense, it is not necessarily the causal interpretation relevant to a specific clinical question. This is particularly important when making treatment recommendations at an individual level, where clinicians and patients may be more interested in understanding how a given treatment would affect a specific patient rather than an average comparison across populations.

### v3

# Interpretational challenges of the Win Ratio in analyzing Hierarchical Composite Endpoints in Chronic Kidney Disease

@@ Abstract
Win statistics based methods have gained traction as a method for analyzing Hierarchical Composite Endpoints (HCEs) in randomized clinical trials, particularly in cardiovascular and kidney disease research. HCEs offer several key advantages, including increased statistical power, mitigation of competing risks, and hierarchical ranking of clinical outcomes. While, as summary measures, the win ratio (WR) along with the Net Benefit (NB) and the Win Odds (WO) provide a structured approach to analyzing HCEs, several concerns regarding their interpretability remain. In this paper, we present known issues with the WR using simple examples designed to explore the implications for the clinical interpretability of the treatment effect measure in the chronic kidney disease setting. Specifically, we discuss the challenge of defining an appropriate estimand in the context of HCEs using the WR, the difficulties in formulating a relevant causal question underlying the WR, and the dependency of the WR on the variance of its components, which complicates its role as an effect measure. Additionally, we highlight the non-collapsibility and non-transitivity of the WR, further complicating its interpretation. While the WR remains a valuable tool in clinical trials, its inherent limitations must be acknowledged to ensure its proper use in regulatory and clinical decision-making.
Keywords: Win Statistics, Estimand.
@@ 1 Introduction
In recent years, there has been a notable surge in the analysis of Hierarchical Composite Endpoints (HCE) using the Win Ratio (WR) in randomized clinical trials. Mao’s study [1], based on a search of ClinicalTrials.gov, highlights a significant rise in the application of WR analyses. Within the realm of cardiovascular and kidney disease, several major outcome trials have employed the WR in some capacity to compare a novel treatment intervention with a comparator [2, 3, 4, 5, 6, 7, 8].
The application of the Win Ratio (WR) in clinical trials has drawn critiques from a clinical perspective; see, for example, [9] and the discussion by [10, 11]. Our aim in this article is to present already established statistical issues with the WR from a perspective of clinical interpretability, using simple examples designed to enhance understanding. For a comprehensive overview of the general pairwise comparison approach to data analysis, consult [12]. As context for our discussion, we focus specifically on two-armed clinical trials evaluating the efficacy of an active treatment against a comparator.
@@ 1.1 Hierarchical composite endpoint
A common feature of the HCEs used in almost all of the above-mentioned trials is that they typically involve one terminal event, one or more non-terminal events, and one or more continuous/ordinal outcomes. For example, [2], defined a heart failure HCE as "...a hierarchical composite of death, number of HF events, time to first HF event, or a $\geq$ 5-point difference in Kansas City Cardiomyopathy Questionnaire (KCCQ) total symptom score change at 90 days." Similarly, [4] proposed and investigated a kidney disease HCE comprising "(1) all-cause mortality; (2) kidney replacement therapy defined as dialysis for at least 28 days or kidney transplantation; (3) sustained GFR <15 mL/min per 1.73 $\text{m}^{2}$ for at least 28 days; (4) sustained GFR decline from baseline for at least 28 days of 57%; (5) 50%; (6) 40%; or (7) GFR slope". GFR slope refers to the individual-level, annualized rate of decline in glomerular filtration rate (GFR), which represents kidney function.
@@
While the ordering of the HCE may seem straightforward, it is not always the case. To illustrate this with an extreme example, suppose a trial where all subjects in the comparator arm systematically experience the terminal event slightly earlier than those in the active arm. At the same time, all active arm subjects experience a non-terminal severe event within the HCE, and none in the comparator arm do. In this case, win statistics based on the HCE will indicate superiority for the active arm due to the small benefit on the terminal event, yet the HCE ignores the increased frequency of non-terminal severe events in the active arm. These limitations are also present in other analyses of composite endpoints; for example, a time-to-first-event analysis of this type of data would encounter similar issues. To address these concerns, it is in general advisable to also evaluate the components individually [13, 14]
@@ 1.2 Win Ratio
The WR, as proposed by [15], serves as a summary measure for analyzing HCEs when comparing two treatment arms. In the unmatched version of this approach, each subject from the novel treatment arm is compared to each subject from the comparator arm to determine a "winner" for each comparison. Note that although "success" might be considered a more appropriate term in some contexts, the use of "winner" is well established in the literature. Initially, the analysis determines whether one subject wins over the other on the first HCE component. If a clear winner emerges, that subject is declared the winner for that specific subject comparison. If a winner cannot be determined, a tie is declared for this component, and the process moves on to the next component. This sequence continues until a winner is declared for the entire comparison or until a tie is declared for the last component, resulting in an overall tie for that particular subject comparison. Let $P_{t}$ denote the proportion of pairs when treatment wins, $P_{c}$ denote the proportion of pairs when control wins, and $P_{\text{tie}}=1-P_{t}-P_{c}$ denote the proportion of ties. The WR is then estimated as $P_{t}/P_{c}$. Consequently, the win ratio does not account for ties, and in situations where the proportion of ties is large, it may provide a misleading measure, similar to the risk ratio in analyses of dichotomous endpoints when absolute risks are small. In addition to concerns about ties, another important consideration in the application of the WR is how to account for baseline covariates that may influence outcomes. Covariate adjustment of the win ratio is an area in development, see e.g. [16]; a simple, often-employed method is stratification. The stratified win ratio considers predefined strata and is calculated by comparing the number of wins for each treatment group within each stratum, then aggregating these results with appropriate weights.
@@
Other often used summary measures for HCEs are the net benefit (NB) and the Win Odds (WO). The NB, also referred to as the Win Difference, computes the difference in win probabilities between treatment arms, i.e. $P_{t}-P_{c}$. The WO, estimated as $(P_{t}+0.5P_{\text{tie}})/(P_{c}+0.5P_{\text{tie}})$, adjusts for tied outcomes by allocating half of the ties to the wins for each arm, and thus ’penalizing’ trials with many ties. While these three win statistics differ in their interpretations and sensitivity to ties, their corresponding statistical tests are asymptotically equivalent under standard regularity conditions [17, 18]. We will focus on the WR in this article, but briefly mention NB and WO when relevant.
@@
All of the benefits and issues presented in this section are to be considered when applying a WR to HCEs. However, in this article, we will direct our focus toward specific aspects related to the interpretation of the WR estimate. First, we will delve into the challenge of defining a suitable estimand for which a WR analysis of an HCE would provide a proper estimator, particularly formulating the clinical question that the WR would aptly address. Second, we will tackle the difficulty of defining a relevant causal question that can be answered by a WR analysis. Third, we also discuss the non-transitivity of the WR. Fourth, we will address the fact that a WR of an HCE with one or more continuous components will be contingent on the variance of those components. As a consequence, the WR is also a non-collapsible measure (similar to hazard or odds ratios). Taken together, these considerations add significant complexity to the interpretation of the WR as an effect measure
@@
All these factors collectively lead to the acknowledgment that the win ratio is problematic as an effect measure, warranting further methodological refinement.
@@ 2 Estimand
The estimand framework was introduced in 2020 with the publication of the ICH E9 addendum [19]. This framework provides a structured approach to defining estimands in clinical trials, ensuring clarity, transparency, and interpretability of clinical trial results, thereby improving decision-making processes in drug development and regulatory assessments. The key attributes of the estimand include:
@@
  • Treatment: Assigned novel or alternative treatment, on top of standard of care.
@@
  • Population-level Summary: The ratio of the probability that a random subject assigned to the intervention has a better outcome to the probability that a random subject assigned to intervention has a worse outcome.
@@
A fundamental consideration when defining an estimand for a WR-based analysis is its dependency on the maximum follow-up time and the censoring distribution. Since the WR is based on pairwise comparisons between subjects, its estimation is affected by the time window over which these comparisons are conducted  [20]. To define a valid estimand that is independent of censoring, one approach is to restrict WR calculations to a fixed time horizon $t$, such as 24 months [21, 22], ensuring that any observed differences in follow-up arise exclusively from loss-to-follow-up. This reduces interpretational complexities from differential follow-up, as the contributions of the more severe endpoints comprising the HCE tend to increase with longer follow-up time; thus, the relative contributions of the different endpoints will generally vary between patients with shorter versus longer follow-up durations. It also improves comparability across studies, provided that the fixed follow-up period is identical in each study
Alternatively, if clinically justified, the WR can be assumed to remain constant over time, allowing estimation of a single value representing the entire study period [23, 24]. Despite these considerations, testing H0 : WR = 1 remains valid, even if its interpretation depends on follow-up duration. The usual statistical tests for WR, such as asymptotic methods or permutation-based approaches, remain robust for evaluating the null hypothesis of no treatment effect. However, careful interpretation of the WR estimate is necessary to ensure it aligns with the clinical question of interest.
@@
If the primary clinical interest lies in comparing the location of a continuous endpoint between treatment arms, in a population where a significant number of deaths is anticipated during the trial, one might consider employing an HCE with time to death as the first component, and the continuous endpoint as the second in order to circumvent interpretational challenges resulting from early termination of follow-up by "truncation by death", thus, in this context, the clinical event of death act as nuisances that complicate the evaluation of the continuous endpoint. However, if we are to remain faithful to the strict estimand framework, the technical challenges of managing intercurrent events, such as death, in the comparison of continuous outcomes should not dictate the choice of the estimand. Specifically, if the estimand aims to assess the treatment effect on the continuous endpoint, the composite strategy of handling intercurrent events through an HCE will not adequately address the estimand of interest. Moreover, if applied to the normally or log-normally distributed continuous endpoint, the WR as a treatment effect measure does not estimate the difference in the locations of the continuous distributions (as can be seen from the example below of normally distributed random variables, WR is in this case related to the standardized mean difference, see equation (1) and not the mean difference); rather, it assesses the odds that a randomly selected subject assigned to the intervention has a better outcome than a randomly selected subject assigned to the control group.
On the other hand, in the CKD example the primary clinical interest lies in the treatment effect on kidney function decline. Thus endpoint components of interest would be GFR slope and relevant clinical events such as dialysis or kidney transplantation. In this setting, these clinical events are essential and intentionally incorporated into the composite endpoint, as they provide crucial insights into kidney function. Interpreting GFR values after dialysis or transplantation is not straightforward. Hence, the estimation of GFR slope in the presence of these intercurrent events, which are by themselves highly relevant for the evaluation of kidney function, may also be challenging. To overcome these issues, employing the HCE for the estimand as outlined in the CKD example above might be an attempt for a holistic assessment of kidney function.
@@ 2.1 Causal interpretation
As mentioned in section 2, the ICH E9 addendum on estimands emphasizes the importance of clearly defining the treatment effect of interest in clinical trials. The addendum encourages the use of causal language and frameworks to describe these effects, ensuring that the treatment effect is interpretable in a causal manner. For an introductory text on causal inference please refer to [25].
@@
where comparisons are done within subjects. However this is not identifiable in a randomized experiment [26, 27].
There is no guarantee that these two are similar. One can even construct examples with opposing conclusions. As an illustration, consult the so-called Hands paradox [28, 26]: in Table 1, we present the factual and counterfactual responses for three subjects under two treatments. Here, Y(0) represents the response under the comparator treatment, while Y(1) denotes the response under the active treatment. A higher response value indicates a better outcome. In this toy example, the WR (population-level), comparing all outcomes in the two arms is: ((0+0+0) + (0+1+0) + (0+1+1)) / ((1+1+1) + (1+0+1) + (1+0+0)) = 3/6 = 0.5, whereas the $\widetilde{\text{WR}}$ (individual-level), doing the comparisons within individuals (0+1+1)/(1+0+0) = 2/1 = 2. Thus, not only are they different, but they are pointing in opposite directions; one would lead to a conclusion of a beneficial individual-level treatment effect, whereas the population-level would suggest a harmful treatment effect.
@@
The non-identifiable individual-level $\widetilde{\text{WR}}$ has a straightforward interpretation for patients and physicians: "What is the ratio of the probability of obtaining a better outcome to the probability of a worse outcome when comparing treatment 1 with treatment 0". However, the interpretation of the population-level WR for a subject/physician is more dubious: "what is the ratio of the probability of obtaining a better outcome to the probability of a worse outcome for a subject exposed to treatment 1 vs another subject exposed to treatment 0". While this population-level comparison is valid in a causal sense, it is not necessarily the causal interpretation relevant to a specific clinical question. This is particularly important when making treatment recommendations at an individual level, where clinicians and patients may be more interested in understanding how a given treatment would affect a specific patient rather than an average comparison across populations. The same interpretational concerns apply to the WO and NB, as they are fundamentally based on the same win probability comparisons as the WR. A similar exposition is given in [29]. This distinction between individual-level and population-level causal interpretations has been explored in other works as well, including a detailed exposition by [29].
3 Non-transitivity
A natural expectation from any effect measure is that it adheres to the principle of transitivity. Transitivity implies that if treatment A is deemed superior to treatment B, and treatment B is deemed superior to treatment C, then one should logically conclude that treatment A is superior to treatment C. This property provides coherence and consistency in comparative analyses. Lack of this property would render the effect measure a possibly unreliable tool for decision-making, as it not only complicates clinical decision-making but also limit the effectiveness of meta-analytic approaches, where cross-trial consistency is critical.
Non-transitivity may arise in WR, WO, and NB calculations due to censoring of time-to-event components. This issue occurs when subjects are observed over varying time intervals, necessitating comparisons based on overlapping observation periods. For further details, see [20], and [30].
Non-transitivity can also occur for continuous or ordinal components. For example, consider three treatments (A, B, and C) with outcomes shown in Table 2. When comparing treatment pairs using the WR, we observe that WR(A vs B) = WR(B vs C) = WR(C vs A) = $1.25$, which violates transitivity, in other words, no single treatment emerges as the best because the treatments each outperform each other in a circular manner. This example corresponds to a simplified version of Efron’s dice. A similar argument for the Wilcoxon–Mann–Whitney test is given in [31].
Interestingly, non-transitivity persists even when considering the non-identifiable individual-level effect measure $\widetilde{\text{WR}}$, as illustrated by the example in Table 2. When the data in the table is interpreted as factual and counterfactual outcomes for three hypothetical subjects, the calculated $\widetilde{\text{WR}}$(A vs B) = $\widetilde{\text{WR}}$(B vs C) = $\widetilde{\text{WR}}$(C vs A) = $2$. Importantly, the challenges discussed here are not exclusive to the WR but extend to the WO and NB as well.
Table 2.
Subject A B C
1 2 1 3
2 4 6 5
3 9 8 7
Table 2: Toy example with outcomes across three treatment arms.